OpenStudy (jake124):

you mix the letters M A T H M A T I C A l thoroughly without looking you draw one letter find the probabilaty write the probabilityas a fraction in simplest form a decimal please help a percent

OpenStudy (jake124):

@lilkg77

OpenStudy (sshayer):

count the total number of letters

OpenStudy (jake124):

11

OpenStudy (campbell_st):

just use P(A) = number of A's/total number of letters

OpenStudy (sshayer):

one by one choose number of M's

OpenStudy (jake124):

im confused

OpenStudy (jtvatsim):

Are we missing something in the question? It seems like the question got cut off. It says "find the probability" but what are we finding the probability of?

OpenStudy (jake124):

no i wrote the question how it was wrote

OpenStudy (sshayer):

\[P(M)=\frac{ number~ of ~M's }{ total~ number~ of~ letters }\]

OpenStudy (campbell_st):

so you need to find the probability of an A being selected is that the way the question reads

OpenStudy (jake124):

yes

OpenStudy (sshayer):

similarly for other letters

OpenStudy (jake124):

whats the answer

OpenStudy (sshayer):

you cont number of M's

OpenStudy (jake124):

theres 2 m's

OpenStudy (sshayer):

so probability of drawing M will be=?

OpenStudy (jake124):

2 ?

OpenStudy (sshayer):

formula is above written

OpenStudy (sshayer):

you have to fill the values only.

OpenStudy (jake124):

im sorry im not really good at math can you explain ?

OpenStudy (sshayer):

\[P(M)=\frac{ 2 }{ 11 }\]

OpenStudy (sshayer):

similarly you show me for A

OpenStudy (jake124):

11 ?

OpenStudy (sshayer):

\[P(A)=\frac{ number ~of A's }{ total ~number~of~letters }\]

OpenStudy (sshayer):

How many A's?

OpenStudy (jake124):

there a 3 As

OpenStudy (sshayer):

then write the probability of A

OpenStudy (jake124):

3

OpenStudy (sshayer):

how i have found the probability of M similarly P(A)=?

OpenStudy (jake124):

so 3 is right

OpenStudy (sshayer):

3 is not probability it is number of A's

OpenStudy (jake124):

so find the probaility of A

OpenStudy (sshayer):

formula for P(A) is written above.

OpenStudy (jake124):

3/4 ?

OpenStudy (sshayer):

from where you get 4.

OpenStudy (sshayer):

total number of letters=11

OpenStudy (jake124):

im running out of time i nedd the answer

OpenStudy (sshayer):

\[P(A)=\frac{ number~of~A's }{ total~number~of~letters }=?\]

OpenStudy (sshayer):

\[P(A)=\frac{ 3 }{ ? }\]